# Don’t Pay Attention

### v1

@@ Appendix I Learning Rate: To Decay or Not to Decay?
To this end, we adopt a cosine decay schedule with a peak learning rate of $1e{-3}$ for Avey, especially that it provides the lowest loss across all the runs. We note, however, that Table 10 also shows that the lowest loss does not correspond to the best downstream task performance. For instance, the loss of 3.308 under the constant learning rate schedule resulted in a slightly better benchmark performance than the lower loss of 3.218 under the cosine decay schedule, both using the same peak learning rate. While constant learning rates can be effective for short or exploratory runs (this study uses only 10 billion tokens), it is generally the case that, as the number of training tokens increases, the learning rate must decrease to allow the optimizer to settle into a lower-loss region [103]. Hence, schedules with decay are typically favored for longer or large-scale training runs [35, 4].
Appendix J Scaling Laws
@@
Mamba-153M 28 768 1.00e-03 2B
Mamba-496M 42 1280 1.00e-03 7B
Mamba-1.5B 52 2048 1.00e-03 20B
RWKV-7-152M 12 768 3.00e-03 2B
RWKV-7-488M 30 1024 1.50e-03 7B
RWKV-7-1.5B 24 2048 1.25e-03 20B
@@
In this section, we present a scaling law study comparing how well Avey, Transformer++, Mamba, and RWKV-7 scale with increasing compute. For Avey, we use the full architecture, including both the ranker and neural processor. All models are trained at three different sizes, as defined in Appendix A and summarized in Table 11. To ensure compute-optimal scaling, we proportionally increase the number of training tokens with model size, following the Chinchilla scaling laws [35]. Specifically, and in line with the methodology from [25], we use 2B, 7B, and 20B tokens to train models with approximately 150M, 500M, and 1.5B parameters, respectively. Lastly, we employ the same batch size across all the models to control for variability in the number of gradient update steps, especially because of training with a limited number of tokens.
@@
Fig. 6 presents the scaling results. The x-axis represents the total training compute budget, calculated as the product of the number of training tokens and model parameters, which serves as a proxy for the total FLOPs required to train each model. As shown, Avey exhibits the steepest decline in perplexity as compute increases. Although it begins with a relatively high perplexity [^10] Avey can achieve substantially lower perplexity under alternative configurations. For example, the small model (150M parameters) with sequence length $N=1024$, split size $S=16$, and top-$k=15$ achieves much lower perplexity as shown in Table 8. The configuration used in this experiment– and in Sections 3.3 and 3.4, was selected based on its strong downstream benchmark performance, rather than optimal perplexity. As discussed in Appendix B, perplexity serves as a useful proxy for modeling capability, but does not always align perfectly with downstream task accuracy., it improves more rapidly than the other models, indicating strong scaling behavior and greater benefit from additional compute. Following Avey, Transformer++ demonstrates the next-best scaling trend, outpacing Mamba and RWKV-7. While Mamba achieves relatively low perplexity at smaller compute budgets, it does not scale as effectively as Avey or Transformer++. Finally, RWKV-7 performs well at low compute but shows the flattest scaling curve, suggesting it gains the least from additional training compute.
Appendix K Ablation Study
Table 12: Ablation results comparing Avey variants, with individual components removed or replaced.
Model Variant ppl ARC-C ARC-E Hella OBQA PIQA SIQA Wino Average
In this study, we perform a series of ablation experiments on the core components of Avey, leveraging the best configurations identified in Appendices BCDEFH, and I. All experiments are conducted using Avey’s complete architecture, comprising both the ranker and neural processor, with the small model variant (153 million parameters) as the baseline. For this study, we trained this model on 10 billion tokens from the FineWeb dataset, using the training methodology and hyperparameters detailed in Appendix A.
We begin by examining the effect of dynamic parameterization on both perplexity and downstream benchmark performance. As described in Section 2.2.2, dynamic parameterization allows each neuron in the contextualizer network to compute a cosine similarity between its input embedding and the embeddings of all other neurons, weight those embeddings accordingly, and aggregate them via a learned weighted sum. This mechanism induces selectivity, as defined in [25], into the neural processor, thereby making its parameterization dynamic (or input-dependent). Table 12 summarizes the results. Disabling this component results in a 14.3% increase in perplexity and a 0.8% drop in average performance, highlighting its importance.
Second, we evaluate the impact of the partial-embedding bypassing technique introduced in Section 2.2.1. This method involves forwarding a portion of each expanded embedding directly to the fuser, allowing raw, distinctive features to be preserved and potentially serve in promoting more diverse representations. As shown in Table 12, removing this mechanism results in an 8.5% increase in perplexity and a 2.2% drop in average performance, underscoring its significance.
Third, we set the expansion factor in the enricher to 1, effectively disabling the expansion of input embeddings. As illustrated in Table 12, this modification results in a 33.1% increase in perplexity and a 5.2% drop in average performance, corroborating the critical role of embedding expansion in the model’s effectiveness.
Fourth, we remove the weighting of each selected split by its corresponding normalized MaxSim score, thereby preventing the contextualizer from scaling each split’s contribution during contextualization. As depicted in Table 12, this adjustment leads to a 3.8% increase in perplexity and a 1.37% drop in average performance, indicating the importance of this technique.
Fifth, we evaluate Avey without the ranker to assess its impact on downstream task performance, beyond its primary role of enabling effective extrapolation past the trained context window. As shown in Table 12, the ranker does indeed enhance the neural processor’s performance, primarily by improving the quality of contextualization through more meaningful cross-token interactions. We note, however, the slight increase in loss (by 0.5%), which again highlights (as in Appendix B) the discrepancy between the objectives of pertaining and downstream tasks.
Finally, we replace Avey’s neural processor with self-attention to assess the relative contribution of each component to Avey’s overall performance, given that both are designed to pursue cross-token interactions. As illustrated in Table 12, this substitution leads to a 4.6% increase in perplexity and a 2.1% decline in average performance, underscoring the significance of the neural processor and suggesting that self-attention is less effective within Avey’s architectural framework.
Appendix L Additional Short-Range Benchmark Results

### v2

@@ 3.2.8 Learning Rate: To Decay or Not to Decay?
To this end, we adopt a cosine-decay schedule with a peak learning rate of 1e31e{-3} for Avey, as it yields the lowest training loss across all the runs. We note, however, that Table 10 also shows a mismatch between loss and downstream accuracy. For example, a higher loss of $3.308$ under the constant schedule produced slightly better benchmark performance than a lower loss of $3.218$ under cosine decay at the same peak rate. While constant learning rates can be effective for short or exploratory runs (this study uses 10B tokens), it is generally the case that, as the number of training tokens increases, the learning rate should decrease to allow the optimizer to settle into a lower-loss region (you2019does). Consequently, schedules with decay are typically preferred for longer or large-scale training (hoffmann2022training; bergsma2025straight).
3.3 Ablation Study
Table 11: Ablation results comparing Avey variants, with individual components removed or substituted.
Model Variant Perplexity ARC-C ARC-E Hella OBQA PIQA SIQA Wino Average
In this section, we conduct a series of ablations on Avey’s core components, building on the best configurations identified in Section 3.2. All experiments use Avey’s full architecture, comprising the neural processor and ranker, with the small model (153M parameters) as the baseline. We train this model on 10B tokens from the FineWeb dataset, following the training methodology and hyperparameters detailed in Section 3.1.
We begin by examining the effect of dynamic parameterization on both perplexity and downstream benchmark performance. As described in Section 2.2.2, dynamic parameterization enables each neuron in the contextualizer to compute a cosine similarity between its input embedding and the embeddings associated with other neurons, weight those embeddings accordingly, and aggregate them via a learned weighted sum. This mechanism induces selectivity (gu2023mamba) in the neural processor, making its parameterization input-dependent. Table 11 summarizes the results. Disabling this component increases perplexity by $14.3\%$ and reduces average downstream performance by $0.8\%$, highlighting its importance.
Second, we assess the impact of partial-embedding bypassing introduced in Section 2.2.1. This mechanism forwards a fraction of each expanded embedding directly to the fuser, preserving raw, distinctive features that can promote more diverse representations. As reported in Table 11, removing this component increases perplexity by $8.5\%$ and reduces average downstream performance by $2.2\%$, underscoring its significance.
Third, we set the enricher’s expansion factor to $1$, effectively disabling embedding expansion. As illustrated in Table 11, this change increases perplexity by $33.1\%$ and reduces average downstream performance by $5.2\%$, corroborating the critical role of embedding expansion in the model’s effectiveness.
Fourth, we remove the weighting of each selected split by its corresponding normalized MaxSim score, thereby preventing the contextualizer from scaling each split’s contribution during contextualization. As depicted in Table 11, this adjustment increases perplexity by $3.8\%$ and reduces average downstream performance by $1.37\%$, indicating the importance of this technique.
Fifth, we evaluate Avey without the ranker to assess its impact on downstream performance beyond its primary role of enabling effective extrapolation outside the trained context window. As shown in Table 11, the ranker improves the neural processor’s performance, primarily by enhancing contextualization through more meaningful cross-token interactions. We note, however, a slight increase in loss ($0.5\%$), which again highlights (as in Section 3.2.1) the discrepancy between the objectives of pertaining and downstream tasks.
Finally, we replace Avey’s neural processor with a self-attention module to assess the relative contribution of each component to overall performance, given that both are designed to facilitate cross-token interactions. As illustrated in Table 11, this substitution increases perplexity by $4.6\%$ and reduces average downstream performance by $2.1\%$, underscoring the centrality of the neural processor and suggesting that self-attention is less effective within Avey’s architectural framework.
3.4 Short-Range Benchmark Results
Table 12: Zero-shot performance across multiple NLP tasks.
Model ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
Avey-153M 24.37 42.33 39.36 31.40 68.37 39.13 51.28 42.32
Transformer++-152M 23.63 43.17 39.32 29.80 67.01 38.89 50.22 41.72
Mamba-144M 24.17 43.53 40.55 30.40 68.32 39.41 52.72 42.73
RWKV-7-168M 24.17 43.01 41.55 29.67 68.72 39.17 51.09 42.48
Avey-496M 27.50 48.95 51.82 32.47 72.49 40.15 54.38 46.82
Transformer++-488M 26.73 48.09 52.66 31.73 72.13 39.93 55.25 46.65
Mamba-500M 28.64 51.02 54.15 34.47 73.03 40.84 55.49 48.23
RWKV-7-501M 27.13 49.37 54.54 36.27 73.58 39.40 55.72 48.00
Avey-1.52B 31.26 56.55 61.42 36.80 75.61 42.00 57.06 51.53
Transformer++-1.5B 30.00 56.29 63.87 38.00 76.01 42.24 61.38 52.54
Mamba-1.4B 32.13 57.74 63.74 36.85 76.19 42.00 60.40 52.72
RWKV-7-1.5B 32.94 59.05 64.43 37.13 76.84 41.71 60.06 53.17
In this section, we evaluate Avey on standard autoregressive language-modeling benchmarks and compare it against Transformer++, Mamba, and RWKV-7 across three parameter scales, small, medium, and large (see Table 1). We adopt a suite of widely used zero-shot downstream tasks, as outlined in Section 3.1. Results are summarized in Table 12. For the small models, Avey, Mamba, and RWKV-7 outperform Transformer++ by average margins of $1.43\%$, $2.41\%$, and $1.82\%$, respectively. Additionally, Mamba and RWKV-7 slightly exceed Avey by $0.90\%$ and $0.30\%$. For the medium models, Avey, Mamba, and RWKV-7 again surpass Transformer++ by $0.30\%$, $3.40\%$, and $2.90\%$, respectively. For the large models, Avey underperforms Transformer++ by $1.90\%$, while Mamba and RWKV-7 edge out Avey by $0.71\%$ and $1.19\%$, respectively.
The results above assume a fixed training budget of 100B tokens. To analyze Avey’s scaling behavior with respect to model size, we conduct additional experiments following the Chinchilla scaling laws (hoffmann2022training), which recommend increasing the number of training tokens in proportion to model size. Accordingly, we adjust the training steps and token budgets to align with these prescriptions. All results are reported in Appendix A, which also details the model configurations (e.g., number of layers and embedding dimensions), optimization settings (e.g., learning rates), training schedules, and total tokens for all evaluated models. Our methodology largely follows gu2023mamba, with minor modifications to satisfy parameter-budget constraints. Results in Appendix A show that Avey exhibits the strongest scaling behavior, followed by Transformer++, Mamba, and RWKV-7, in that order.
3.5 Long-Range Benchmark Results
Figure 5: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens). All models use 0.5B parameters. Similar results are shown in Appendix C for other model sizes.
We now evaluate Avey, Transformer++, Mamba, and RWKV-7 on benchmarks targeting long-range dependencies. Specifically, we use the standard Single Needle-In-A-Haystack (S-NIAH) suite from RULER (hsieh2024ruler), as described in Section 3.1. The suite includes multiple variants, notably S-NIAH-1 and S-NIAH-2. S-NIAH-1 requires retrieving the value associated with a given key (the pass-key) from a distractor context (the haystack) containing many key–value pairs. The pass-key (or the needle) appears exactly once, and the model must recover its corresponding value regardless of position within the haystack. S-NIAH-2 follows the same structure but is more challenging, whereby the target value is numeric (e.g., a random 9-digit number), demanding exact recall in which any single-digit error is scored as incorrect. Consequently, this variant tests precise extraction of structured numerical information from long contexts.
Fig. 5 demonstrates the results for Avey, Transformer++, Mamba, and RWKV-7 on both S-NIAH-1 and S-NIAH-2. As noted in Section 3.1, Transformer++, Mamba, and RWKV-7 were all trained with a 2,048-token context window. Transformer++ performs strongly on both benchmarks while the haystack length remains within this window. Once the haystack length exceeds the window width, the accuracy of Transformer++ collapses. Mamba and RWKV-7 generalize somewhat beyond their training windows, but their performance also degrades substantially as the haystack grows. In contrast, Avey maintains high accuracy on both benchmarks despite being trained with a 512-token context window. For example, on S-NIAH-2 with a 64k-token haystack, Avey outperforms Mamba and RWKV-7 by average margins of $85.25\%$ and $23.6\%$, respectively. On S-NIAH-1 under the same 64k-token setting, Avey attains $97.8\%$ accuracy, whereas Mamba and RWKV-7 drop to $0\%$ and $0.8\%$, respectively.
Interestingly, Avey’s performance often improves as the haystack length increases, highlighting strong extrapolative capability. One plausible explanation is that as the haystack (or the sequence length $N$) grows, the candidate pool from which the ranker selects top-$k$ splits also expands. As discussed in Section 3.2.6, a larger $N$ enables the ranker to retrieve more relevant splits while discarding less informative ones, improving the quality of contextualization and potentially boosting accuracy. This interpretation is consistent with findings in Section 3.3, where introducing the ranker yielded measurable gains. Notably, embeddings containing a needle (whether in S-NIAH-1 or S-NIAH-2) are not processed in isolation but are contextualized alongside other embeddings. As such, better ranker-driven context selection can result in more precise value recall. That said, whether this fully explains Avey’s increasing accuracy with longer haystacks remains an open question, and further interpretability studies are needed to better understand the underlying drivers of this behavior.
4 Related Work
Recurrent Neural Networks (RNNs) (elman1990finding; rumelhart1986learning) are designed to process sequential data by capturing temporal dependencies, making them well-suited for tasks where input order is essential. However, their cyclical nature limits their potential for parallel computation and exposes them to vanishing and exploding gradient problems. As a result, they typically struggle to effectively learn long-range dependencies. While architectures like Long Short-Term Memory (LSTM) (hochreiter1997long) and Gated Recurrent Units (GRU) (cho2014learning) mitigate these gradient-related issues, they remain slow to optimize and challenging to scale due to retaining RNN’s core recurrent and sequential structure.
In contrast, the Transformer (vaswani2017attention) employs a self-attention mechanism to process each sequence of tokens simultaneously. More precisely, it promotes two key design principles: (1) a recurrent-free architecture, which enables parallel computation of token embeddings, while still capturing their order through positional encodings, and (2) a multi-head self-attention approach, which facilitates cross-token interactions to further enrich the expressiveness of embeddings. These innovations make the Transformer highly effective, as well as parallelizable and efficient to train. However, they also cause its computational and memory requirements to scale quadratically with sequence length, making it expensive and less efficient for very long sequences.
To address the Transformer’s quadratic computation and memory costs, a wide array of approaches have been proposed, including linear attention (kitaev2020reformer; katharopoulos2020transformers; choromanski2020rethinking; peng2021random; zhai2021attention), sparse or local attention (yuan2025native; child2019generating; parmar2018image), context compression (rae2019compressive; wang2020linformer; sukhbaatar2019adaptive; roy2021efficient), and modified attention computations (tay2021synthesizer; wu2019pay; tay2020sparse), to mention just a few. Notably, the Attention-Free Transformer (AFT) (zhai2021attention) offers a linear drop-in replacement for the quadratic self-attention mechanism. In particular, it weights key and value vectors using learned positional biases and integrates them with query vectors via element-wise multiplication. As such, it eliminates the need to compute and store the costly attention matrix while still preserving global query-value interactions—without requiring architectural modifications or additional tuning. Furthermore, AFT introduces variants such as AFT-local and AFT-conv, which leverage local attention patterns to reduce parameter count and further improve computational and memory efficiency.
RWKV-4 (peng2023rwkv) (the first 3 versions were experimental (li2024survey)) capitalizes on AFT and suggests combining the strengths of both Transformers and RNNs. To elaborate, unlike Transformers and akin to RNNs, it does not process each input token solely based on its own embedding, but rather as a weighted sum of its embedding and that of the preceding one. To the contrary of traditional RNNs and similar to Transformers, it adopts self-attention, but an extended version of it, namely, that of AFT. This hybrid approach allows RWKV-4 to maintain some of the modeling capabilities of RNNs, while benefiting from the parallelization and scalability features of Transformers.
More precisely, RWKV-4 extends AFT in two distinct ways: (1) it introduces an additional parameter to handle each current token independently, and (2) it implements a per-time-step decay mechanism that selectively removes older content from the recurrent hidden state in a data-dependent manner. This decay mechanism addresses a central limitation of linear attention, which pertains to the lack of a systematic way to discard outdated information (schlag2021linear; yang2023gated).
RWKV-5 (peng2024eagle) enhances RWKV-4’s architecture and learning decay mechanism by replacing traditional vector-valued states with more expressive multi-head, matrix-valued ones. Moreover, it reconfigures receptive states, incorporates supplementary gating mechanisms, and dynamically learns the linear interpolation between the current and preceding token embeddings instead of relying on pre-defined hyperparameters.
RWKV-6 (peng2024eagle) promotes a new application of low-rank adaptation functions (hu2022lora; li2024survey). Specifically, it makes the linear interpolation between the current and preceding tokens data-dependent to improve the selectivity of the model in retaining and discarding information. Additionally, it replaces the static decay mechanism with a dynamic one, allowing each element in the decay vector to fluctuate independently over time in response to the input.
The decay strategies of RWKV-4, RWKV-5, and RWKV-6 still cannot remove values stored at specific keys. DeltaNet (schlag2021linear) overcomes this drawback by partially replacing the values stored at current keys with equivalent new values, enabling models to erase outdated memories and include up-to-date ones on a per-key basis. However, it only allows a fixed scalar fraction of a value to be replaced from a state via an in-context learning rate parameter, thus demonstrating rigidity in adapting to varying data contexts.
RWKV-7 (peng2025rwkv) builds upon the principles of DeltaNet and introduces a vector-valued in-context learning rate instead of a scalar-valued one. This allows selective replacement of state data on a channel-wise basis. Furthermore, RWKV-7 employs a vector-valued decay mechanism and uses additional low-rank projections to optimize the trade-off between the number of parameters, computational efficiency, and downstream performance. Lastly, it incorporates Value Residual Learning (zhou2024value), which improves the propagation of initial local information via utilizing a residual connection between the value vectors of the current layer and those of the first layer prior to the attention operation, resulting in enhanced language modeling performance.
Most recently, RWKV-X (hou2025rwkv) proposed combining the strengths of RWKV and sparse attention, drawing inspiration from Mixture of Block Attention (MoBA) (lu2025moba). In particular, RWKV-X restricts each query to attend only to a small, relevant subset of the input, thus reducing computational cost and facilitating the modeling of longer-range dependencies. More precisely, rather than allowing each token to attend to every other token in the sequence (as in traditional self-attention), it constrains each token’s attention to a limited subset (hence, making it sparse), while maintaining the coupling between the input sequence and context window.
Similar to RWKV, RetNet (sun2023retentive) adopts linear attention and promotes a hybrid approach that blends Transformer- and RNN-like representations, yet with a decay-based memory unit. Specifically, it divides the input sequence into chunks, wherein the Transformer-like parallel representation is applied. Additionally, it enables propagating information sequentially across chunks using the RNN-like representation. Lastly, it uses multiple attention heads, each governed by a distinct decay rate, and replaces LayerNorm (ba2016layer) with GroupNorm (wu2018group).
Although linear attention has been proposed as a promising alternative to quadratic softmax attention (katharopoulos2020transformers; choromanski2020rethinking; kasai2021finetuning; peng2021random), existing implementations of it are in practice slower than optimized versions of softmax attention (dao2022flashattention; dao2023flashattention; yang2023gated). From an accuracy standpoint, linear attention generally underperforms conventional softmax attention, sometimes by a significant margin in language modeling (kasai2021finetuning; yang2023gated).
To this end, and in light of the exponentially growing complexity associated with overcoming the limitations of Transformer-based architectures, there has been a renewed interest in RNN-based alternatives in recent years. Notably, Structured State Space Sequence (S4) models (gu2021efficiently; gu2021combining), inspired by the classical state space models (SSMs) (kalman1960new), have emerged as a promising paradigm for sequence modeling. These models describe the temporal evolution of a system using differential equations, offering a continuous-time formulation of dynamics, and can be viewed as generalized versions of RNNs.
An SSM as a concept has a broad meaning, which simply refers to the notion of any recurrent process with a latent state (gu2023mamba). From this perspective, the RNN-like linear attention model proposed and formulated by (katharopoulos2020transformers) can be interpreted as a degenerate linear SSM. Interestingly, this justifies the usage of a decay factor in RetNet and RWKV, especially that a decay term (or a forget gate) has been shown to be crucial in RNNs (hochreiter1997long; van2018unreasonable; cho2014learning).
Numerous variants of SSMs (gu2021efficiently; gu2022parameterization; gupta2022diagonal; li2024makes; ma2022mega; orvieto2023resurrecting; smith2022simplified) have demonstrated strong performance across a range of domains, including audio and vision (goel2022s; nguyen2022s4nd; saon2023diagonal). Nonetheless, these variants have struggled with language modeling, often lagging behind Transformers by several points in perplexity (gu2021efficiently).
From an efficiency standpoint, however, SSMs have shown encouraging results in language modeling. For instance, S4 (gu2021efficiently; gu2021combining), a prominent SSM, converts the continuous-time state update equation of SSMs into a discrete form, hence, enabling parallel sequence modeling. Moreover, it utilizes the HiPPO (High-Order Polynomial Projection Operator) initialization (gu2020hippo), which alleviates the vanishing gradient problem and facilitates processing longer sequences.
Another example of SSMs is H3 (fu2022hungry), which improves language modeling by allowing both, the recall of earlier tokens and token-wise comparisons within a sequence. H3 extends S4 by suggesting a state-passing algorithm that enhances computational efficiency on modern hardware. This advancement reduces the hardware-related barriers that have traditionally limited the scalability of SSM-based architectures.
Hyena (poli2023hyena) capitalizes on H3 by replacing its S4 layer with an MLP-parameterized global convolution (romero2102ckconv). S5 (smith2022simplified) proposes using parallel scan (martin2017parallelizing) to parallelize S4. Liquid S4 (hasani2022liquid) augments S4 with an input-dependent state transition matrix, computed convolutionally in the frequency domain (which is computationally efficient) and mapped back to the time domain using an inverse Fourier transformation. SGConv (li2024makes), LongConv (fu2023simple), MultiresConv (shi2023sequence), and Toeplitz Neural Network (qin2023toeplitz) all focus on the convolutional representation of S4 as well, aiming to enhance its efficiency (gu2023mamba).
Most recently, Mamba (gu2023mamba) introduced a new class of SSMs known as selective SSMs, specifically designed to improve the performance of language modeling tasks. Mamba addresses a key limitation in SSMs, namely, their inability to selectively process inputs in an input-dependent manner (i.e., focus on or ignore specific parts of the input sequence). Consequently, it makes the SSM parameters input-dependent, but introduces a technical challenge since traditional SSMs are inherently designed to be time- and input-invariant to ensure computational efficiency. To overcome this challenge, Mamba proposes a hardware-efficient parallel scan (or prefix sum) algorithm (blelloch1990prefix), which enables recurrent-style computation without explicitly materializing the expanded state. This design precludes costly I/O operations across GPU memory hierarchies and accelerates both, training and inference.
Tri Dao and Albert Gu (dao2024transformers) argue that various approaches to operating SSMs can be reframed as matrix multiplication algorithms involving a specific class of structured matrices known as semiseparable matrices. They further leverage the language of tensor contractions to prove the recurrent formulation of linear attention as proposed by (katharopoulos2020transformers), before generalizing it to a new family of structured masked attention (SMA).
Subsequently, Tri Dao and Albert Gu demonstrated that SSMs and Transformers are fundamentally connected, governed by the mathematical framework of semiseparable matrices and SMA. Additionally, they developed a rich state space duality (SSD) framework of theoretical connections between SSMs and various forms of attention. This framework facilitated the design of Mamba-2, an extended version of Mamba, which aims to improve its efficiency (not performance). Mamba-2 utilizes a scalar data-dependent gating mechanism (like the ones proposed by (peng2021random; sun2023retentive; beck2024xlstm)), which enables transforming its recurrent structure into a matrix-multiply form, thus allowing for efficient execution on tensor cores and better support for larger hidden state sizes.
The strategy of the SSD framework mirrors that of linear attention (katharopoulos2020transformers), which established a connection between autoregressive attention mechanisms and linear RNNs via showing an equivalence between "dual forms" of quadratic kernelized attention and a specific type of linear recurrence. Conceptually, the SSD framework seeks to transfer algorithmic and systems-level optimizations originally developed for Transformers to the realm of SSMs. Its overarching goal is to enable the development of architectures that outperform Transformers, while scaling more efficiently with sequence length.
Finally, several works, including Tolstikhin et al. (tolstikhin2021mlp), Melas-Kyriazi (melas2021you), Touvron et al. (touvron2022resmlp), and Ding et al. (ding2021repmlp), among others, have questioned the necessity of self-attention, particularly in the context of Vision Transformers. In contrast, Liu et al. (liu2021pay) introduced gMLP, an MLP-based alternative to BERT-style Transformers (devlin2019bert) that (partially) eliminates self-attention but ultimately underperforms average performance on downstream NLP tasks. gMLP encompasses channel (hidden) and spatial (cross-token) projections with multiplicative gating and static parameterization. Its gating mechanism is reminiscent of Gated Linear Units (GLUs) (dauphin2017language; shazeer2020glu; wu2019pay), as well as earlier architectures such as Highway Networks (srivastava2015highway) and LSTM-RNNs (hochreiter1997long). A key distinction, however, is that gMLP applies gating on the spatially projected dimension and not the hidden one. The gated embedding-wise neural network in Avey’s contextualizer draws inspiration from gMLP.
Unlike all previously mentioned models, Avey abandons self-attention and recurrence, introducing a new architecture composed of a ranker and a dynamically parameterized neural processor. The ranker identifies the most relevant tokens for contextualization, while the neural processor contextualizes them data-dependently. This design decouples sequence length from context width, enabling efficient processing of arbitrarily long sequences without diminishing the influence of distant yet important tokens.
At the core of Avey’s architecture is a weighted-selective-split interaction mechanism, which filters out irrelevant tokens beyond the context window and enables direct interactions only with relevant ones, thus preserving their influence irrespective of sequence length. In addition, Avey employs a partial-embedding bypassing technique that retains a portion of each token’s raw, distinctive features before fusing them with its contextualized ones through a neural network. This technique boosts the performance of Avey (as shown in Appendix 3.3) and might help mitigate issues such as entropy collapse (zhai2023stabilizing) and over-smoothing (zhou2021deepvit; shi2022revisiting), especially at large-scale, when the depth of the model is increased significantly.
5 Conclusion
In this paper, we introduce Avey, a new foundational architecture for autoregressive language modeling. Unlike traditional models, Avey relies on neither recurrence nor attention. Instead, it employs a neural mechanism to enrich and contextualize embeddings, coupled with a ranker that enables flexible and effective handling of arbitrarily long sequences despite being trained with a small context window. We hope this work lays the groundwork for future research and inspires a promising new direction for scalable, effective language modeling.
Acknowledgments
We gratefully acknowledge the Ministry of Communications and Information Technology (MCIT) in Qatar and NVIDIA for providing the computational resources that made many of the experiments reported herein possible. We also thank Microsoft Qatar for technical support on the GPU infrastructure provided by MCIT.
Appendix A Scaling Laws
@@
Mamba-153M 28 768 3.00e-03 2B
Mamba-496M 42 1280 1.50e-03 7B
Mamba-1.5B 52 2048 1.00e-04 20B
RWKV-7-152M 12 768 6.00e-04 2B
RWKV-7-488M 30 1024 4.00e-04 7B
RWKV-7-1.5B 24 2048 4.00e-04 20B
@@
In this section, we present a scaling law study comparing how well Avey, Transformer++, Mamba, and RWKV-7 scale with increasing compute. For Avey, we use the full architecture, including both the ranker and neural processor. All models are trained at three different sizes, as defined in Section 3.1 and summarized in Table 13. To ensure compute-optimal scaling, we proportionally increase the number of training tokens with model size, following the Chinchilla scaling laws [hoffmann2022training]. Specifically, and in line with the methodology from [gu2023mamba], we use 2B, 7B, and 20B tokens to train models with approximately 150M, 500M, and 1.5B parameters, respectively. Lastly, we employ the same batch size across all the models to control for variability in the number of gradient update steps, especially because of training with a limited number of tokens.
@@
Fig. 6 presents the scaling results. The x-axis represents the total training compute budget, calculated as the product of the number of training tokens and model parameters, which serves as a proxy for the total FLOPs required to train each model. As shown, Avey exhibits the steepest decline in perplexity as compute increases. Although it begins with a relatively high perplexity [^12] Avey can achieve substantially lower perplexity under alternative configurations. For example, the small model (150M parameters) with sequence length $N=1024$, split size $S=16$, and top-$k=15$ achieves much lower perplexity as shown in Table 8. The configuration used in this experiment– and in Sections 3.4 and 3.5, was selected based on its strong downstream benchmark performance, rather than optimal perplexity. As discussed in Appendix 3.2.1, perplexity serves as a useful proxy for modeling capability, but does not always align perfectly with downstream task accuracy., it improves more rapidly than the other models, indicating strong scaling behavior and greater benefit from additional compute. Following Avey, Transformer++ demonstrates the next-best scaling trend, outpacing Mamba and RWKV-7. While Mamba achieves relatively low perplexity at smaller compute budgets, it does not scale as effectively as Avey or Transformer++. Finally, RWKV-7 performs well at low compute but shows the flattest scaling curve, suggesting it gains the least from additional training compute.
Appendix B Additional Short-Range Benchmark Results